1/2 points I Previous Answers A spaceship is traveling through deep space towards a space station and needs to make a course correction to go around a nebula. The captain orders the ship to travel 1.9 x 106 kilometers before turning 70° and traveling 1.7 x 106 kilometers before assuming the path towards the space station. If the captain had not ordered a course correction, what would have been the magnitude and direction of the path of the spaceship if it had traveled through the nebula? My Notes Ask Your Teacher space station nebula spaceship magnitude 2.9506 direction e counterclockwise from the +x-axisneed part B please
Answer #1

A=1,9 * 106 i km

-1 . 7 * 106 (cos 70% + sin 70 j ) km

Resultant displacement -A+B=l.9 * 10%+1.7 * 106(cos 70 i + si n 70%)

Ř=(1,9 * 106 + 1.7 * 106 * cos 70) i + 1.7*. 106 sin 70 j

Magnitude of displacement R-1.9106 1.7 106cos70)2(1.7106 sin 70o)2

2,95 106 km

Direction of resultant displacement makes an angle 1.7 106 sin 70 1.9 106 1.7106 cos 70

1.7n 70 1.9+1.7 cos 70° tan1

= tan-1 (1.597477 2.481434 = 32.770 counter clockwise with +x axis.